Which of the following thermodynamic relation is correct ?
(A) dG = VdP – SdT (B) dE = PdV + TdS (C) dH = – VdP + TdS (D) dG = VdP + SdT

Grade:

1 Answers

Vikas TU
3 years ago
Correct ans is dG = dH - Tds - SdT 
as G = H-TS 
H = E+pV 
dH = dE +pdV Vdp 
dG = Vdp - SdT …..................................